# A Linear N-Point Solver for Structure and Motion from Asynchronous Tracks

### v1

@@ 3.3 Solver
One may think that the inversion of $\mathbf{M}_{A}$ in Eq. 8 is expensive since $\mathbf{M}_{A}\in\mathbb{R}^{3M\times 3M}$ leading naively to $O(M^{3})$ complexity. However, the matrix $\mathbf{M}_{A}$ is actually block diagonal with $M$ blocks of size $\times 3$ 3, leading to an efficient inversion algorithm of complexity $O(M)$ instead. Moreover, all terms can be computed from a linear combination of terms $[\mathbf{f}^{\prime}_{ij}]^{2}_{\times}=\mathbf{f}^{\prime}_{ij}{\mathbf{f}^{\prime}_{ij}}^{\intercal}-\|\mathbf{f}^{\prime}_{ij}\|^{2}\mathbf{I}_{3\times 3}$, leading to significant sharing of computation. Finally, having found the estimate $\hat{\mathbf{v}}$ we can find the solution to $\mathbf{P}_{i}$ as
@@ 3.4 Solver Properties
Solution Multiplicity: We start off by discussing the solution multiplicity of the above solver. The SVD operation in Eq. 8 yields two possible unit vectors $\hat{\mathbf{v}}$ and $-\hat{\mathbf{v}}$ which entail the two possible solutions $\hat{\mathbf{P}}_{1:M}$ and $-\hat{\mathbf{P}}_{1:M}$. During deployment, we select the correct solution by recognizing that the recovered points $\hat{\mathbf{P}}_{i}$ must have positive depth. We thus test for the following condition
@@
Degeneracy: Next, let us analyze when degenerate solutions may be encountered. We see that the solution of $\mathbf{v}$ in Eq. 8 depends on the inversion of $\mathbf{M}_{A}$ which in turn depends on the inversion of block diagonal matrices of the form $\mathbf{F}_{i}^{\intercal}\mathbf{F}_{i}$. To successfully invert these matrices, we require that $\mathbf{F}_{i}$ has a full rank. $\mathbf{F}_{i}$ is composed of matrices $[\mathbf{f}^{\prime}_{ij}]_{\times}$ which only have two independent rows. We may thus consider the reduced form of $\mathbf{F}_{i}$ with size $2N_{i}\times 3$ where we have deleted every third row. Thus, to have full rank $N_{i}\geq 2$, i.e. we must enforce that every track has at least two different observations. In practice, we do this by simply discarding tracks with only one observation. Finally, enforcing $\text{rank}(\mathbf{B})\geq 2$ ensures that the SVD step succeeds in Eq. 8. For a discussion on the rank of $\mathbf{B}$ see the appendix.
[…]
@@
  • $M=2$: Here $N\geq 4$. In particular $N_{1}=N_{2}=2$ leads to minimal solver with 888 equations in 888 unknowns. For larger $N$ the system becomes overconstrained again.
  • $M=3$: Here $N\geq 6$ with $N_{1}=N_{2}=N_{3}=2$ yielding a minimal system of 121212 equations in 121212 unknowns.
[…]
@@ 4 Experiments
$\theta_{\text{err}}=\text{arccos}\left(\frac{\mathbf{v}_{\text{gt}}^{\intercal}\hat{\mathbf{v}}}{\|\mathbf{v}_{\text{true}}\|\|\hat{\mathbf{v}}\|}\right)$ (14)
@@ 4.1.1 Analysis of Noise Resilience
[…]
@@ 4.2 Real Experiments
In a next step, we deploy our solver on real data collected with three different sensing modalities (as described in Sec. 3) using public datasets: The Event Camera Dataset [32] was collected using a DAVIS camera [2] which provides a pixel-aligned and time-synchronized stream of global shutter (240×180240$\times$ 180240 × 180@24Hz) and event camera data, as well as IMU measurements. The VECtor dataset [9] provides event camera data and global shutter frames (1224×10241224$\times$ 10241224 × 1024@30Hz) from a stereo rig, as well as IMU data. Finally, the TUM dataset [41] provides rolling shutter images (1280×10241280$\times$ 10241280 × 1024@20Hz). We compare our method against eventail [11], an asynchronous line-based velocity solver. It uses clusters of events, each generated by a separate line to regress velocity components, and fuses these to a full normalized velocity via velocity averaging. On rolling shutter images we implement a similar baseline based on eventail, termed eventail + RS. eventail + RS first extracts Canny edges [3] from the rolling shutter images, and then treats the detected points as events with timestamps assigned according to their row index. We do not report acceleration estimation results on real data due to high noise sensitivity, and leave addressing this challenge for future work.
@@
Table 1: Mean / median velocity error in degrees on tracks from global shutter images. * and gray: results on subset with track inlier ratio >0.9>0.9> 0.9. eventail + E uses events from an event camera.
@@
Table 2: Mean / median velocity error in degrees on tracks from rolling shutter images. * and gray: results on subset with inlier ratio >0.9>0.9> 0.9. eventail + RS uses canny edges from rolling shutter images. “no correction”: no rolling shutter timestamp correction.
@@
Table 3: Mean / median velocity error in degrees on tracks from events. * and gray: results on subset with track inlier ratio >0.9>0.9> 0.9. E stands for events from an event camera. E+GS refers to our method combining tracks from events and global shutter images. Note that for - no collocated GS and event sensor are available.
@@
[…]
@@ Acknowledgments
This research has been supported by project 62250610225 by the Natural Science Foundation of China, as well as projects 22DZ1201900, and dfycbj-1 by the Natural Science Foundation of Shanghai.

### v2

@@ 3.3 Solver
One may think that the inversion of $\mathbf{M}_{A}$ in Eq. 8 is expensive since $\mathbf{M}_{A}\in\mathbb{R}^{3M\times 3M}$ leading naively to $O(M^{3})$ complexity. However, the matrix $\mathbf{M}_{A}$ is actually block diagonal with $M$ blocks of size $3\times 3$, leading to an efficient inversion algorithm of complexity $O(M)$ instead. Moreover, all terms can be computed from a linear combination of terms $\mathbf{f}^{\prime}_{ij}]^{2}_{\times}=\mathbf{f}^{\prime}_{ij}{\mathbf{f}^{\prime}_{ij}}^{\intercal}-\|\mathbf{f}^{\prime}_{ij}\|^{2}\mathbf{I}_{3\times 3}$, leading to significant sharing of computation. Finally, having found the estimate $\hat{\mathbf{v}}$ we can find the solution to $\mathbf{P}_{i}$ as
@@ 3.4 Solver Properties
Solution Multiplicity: We start off by discussing the solution multiplicity of the above solver. The SVD operation in Eq. 8 yields two possible unit vectors $\hat{\mathbf{v}}$ and $\hat{\mathbf{v}}$ which entail the two possible solutions $\hat{\mathbf{P}}_{1:M}$ and $\hat{\mathbf{P}}_{1:M}$. During deployment, we select the correct solution by recognizing that the recovered points $\hat{\mathbf{P}}_{i}$ must have positive depth. We thus test for the following condition
@@
Degeneracy: Next, let us analyze when degenerate solutions may be encountered. We see that the solution of $\mathbf{v}$ in Eq. 8 depends on the inversion of $\mathbf{M}_{A}$ which in turn depends on the inversion of block diagonal matrices of the form $\mathbf{F}_{i}^{\intercal}\mathbf{F}_{i}$. To successfully invert these matrices, we require that $\mathbf{F}_{i}$ has a full rank. $\mathbf{F}_{i}$ is composed of matrices $\mathbf{f}^{\prime}_{ij}]_{\times}$ which only have two independent rows. We may thus consider the reduced form of $\mathbf{F}_{i}$ with size $2N_{i}\times 3$ where we have deleted every third row. Thus, to have full rank $N_{i}\geq 2$, i.e. we must enforce that every track has at least two different observations. In practice, we do this by simply discarding tracks with only one observation. Finally, enforcing $\text{rank}(\mathbf{B})\geq 2$ ensures that the SVD step succeeds in Eq. 8. For a discussion on the rank of $\mathbf{B}$ see the appendix.
Constraint Analysis: The system in Eq. 6 has $3M+2$ unknowns (number of variables minus one for unobservability of scale), and 3N constraints. However, each $\mathbf{f}^{\prime}_{ij}]_{\times}$ only adds two linearly independent constrains, so the number of constraints is actually $2N$. Thus recovering all unknowns needs
@@
  • $M=2$: Here $N\geq 4$. In particular $N_{1}=N_{2}=2$ leads to minimal solver with $8$ equations in $8$ unknowns. For larger $N$ the system becomes overconstrained again.
  • $M=3$: Here $N\geq 6$ with $N_{1}=N_{2}=N_{3}=2$ yielding a minimal system of $12$ equations in $12$ unknowns.
  • $M>3$: Here $N\geq 2M$ leading to an overconstrained system of $2N$ equations in $3M+2$ unknowns.
@@ 4 Experiments
[…]
@@ 4.1.1 Analysis of Noise Resilience
We first evaluate the solver’s robustness under three noise sources: inaccurate point tracking, temporal misalignment, and orientation drift in camera pose. As illustrated in Fig. 2, experiments vary noise level across practical ranges: pixel noise (0 - 5 pixels), timestamp jitter (0 - 50ms), and rotational perturbation (angular velocity noise of 0 - 30 deg/s), while testing three different settings of observation. The results demonstrate that the solver achieves sub-$5^{\circ}$ error at moderate noise levels (1 pixels, 10 ms jitter, 5 deg/s), validating its feasibility in typical operational scenarios. Notably, timestamp jitter exhibits a nonlinear error escalation, with error rising sharply beyond 15 ms. In contrast, pixel noise induces near-linear error scaling, suggesting the tolerance to common feature-tracking inaccuracies, while rotation perturbation has linear impact on the performance. We also notice more observations can effectively mitigate errors due to pixel noise and timestamp jitter, yet yield limited improvement for rotation-induced errors. Nevertheless, external sensor such as IMU can address this limitation by providing rotation-compensated inputs.
@@ 4.2 Real Experiments
In a next step, we deploy our solver on real data collected with three different sensing modalities (as described in Sec. 3) using public datasets: The Event Camera Dataset [32] was collected using a DAVIS camera [2] which provides a pixel-aligned and time-synchronized stream of global shutter ($240\times 180$@24Hz) and event camera data, as well as IMU measurements. The VECtor dataset [9] provides event camera data and global shutter frames ($1224\times 1024$@30Hz) from a stereo rig, as well as IMU data. Finally, the TUM dataset [41] provides rolling shutter images ($1280\times 1024$@20Hz). We compare our method against eventail [11], an asynchronous line-based velocity solver. It uses clusters of events, each generated by a separate line to regress velocity components, and fuses these to a full normalized velocity via velocity averaging. On rolling shutter images we implement a similar baseline based on eventail, termed eventail + RS. eventail + RS first extracts Canny edges [3] from the rolling shutter images, and then treats the detected points as events with timestamps assigned according to their row index. We do not report acceleration estimation results on real data due to high noise sensitivity, and leave addressing this challenge for future work.
@@
Table 1: Mean / median velocity error in degrees on tracks from global shutter images. * and gray: results on subset with track inlier ratio >0.9>0.9. eventail + E uses events from an event camera.
@@
Table 2: Mean / median velocity error in degrees on tracks from rolling shutter images. * and gray: results on subset with inlier ratio >0.9>0.9. eventail + RS uses canny edges from rolling shutter images. “no correction”: no rolling shutter timestamp correction.
@@
Table 3: Mean / median velocity error in degrees on tracks from events. * and gray: results on subset with track inlier ratio >0.9>0.9. E stands for events from an event camera. E+GS refers to our method combining tracks from events and global shutter images. Note that for - no collocated GS and event sensor are available.
@@
Application to Event-based Cameras Finally, we apply our method to tracks derived from an event-based camera, and show results in Tab. 3. Our point solver running on events alone outperforms eventail by $10-30\%$, and this is again due to the use of point tracks instead of lines. Interestingly, frame-based tracks yield better performance than event-based ones, particularly on the VECtor sequences. This aligns with our simulation findings where higher feature density (from frame cameras’ $5\times$ resolution advantage over event cameras) improves spatial sampling. We show the benefit of combining tracks from different sensing modalities, denoted with E+GS. In particular, the sequences shapes_translation and boxes_translation were recorded with a DAVIS camera [2] which features pixels that simultaneously record events and global shutter images. In this setting, we see that adding images significantly reduces errors. Moreover, comparing to Tab. 1 we also see that adding events improves over the global shutter result, highlighting the complementarity of the sensors. This result highlights the benefit of having an asynchronous point solver that can flexibly incorporate both global shutter and event-based observations.
@@ Acknowledgments
This research has been supported by project 62250610225 by the Natural Science Foundation of China, as well as projects 22DZ1201900, and dfycbj-1 by the Natural Science Foundation of Shanghai and project 225354 of the Swiss National Science Foundation.